OpenStudy (sgadi):

Hi friends, I am new to this place. I have system of DE equations to solve. Can anyone help me? Equations are as follows 'D2x=z*Dy-y*Dz' 'D2y=Dz*x-Dx*z' 'D2z=Dx*y-Dy*x' where D2x = time derivative of x, 2 times and Dx = time derivative of x, 1 time

OpenStudy (anonymous):

Ok... so how have you tried solving this system?

OpenStudy (sgadi):

I tried linearizing ... But I am not able to do so. I am now trying to see the model and change something to simplify.

OpenStudy (anonymous):

Yeah, I'm looking at it and having some difficulty as well.

OpenStudy (anonymous):

So, if you integrate D2x, you get Dx=integral of z dy - integral of y dz. Is that correct?

OpenStudy (sgadi):

'D2x=z*Dy-y*Dz' is same as below \[(d^2x/dt^2)=z*(dy/dt)-y*(dz/dt)\]

OpenStudy (anonymous):

Hmmm... yeah, I'm not sure how to solve it explicitly either.

OpenStudy (sgadi):

:-) I think the solution doesnot exists. Anyway thanks a lot my friend.